OpenStudy (anonymous):

Surface Area of a Regular Pyramid Solve for l: S=B+1/2Pl please show the steps thank you so much

OpenStudy (anonymous):

subtract B\[s-B=\frac{1}{2PI}\]clear denominators (cross multiply)\[2PI(s-B)=1\]divide by 2P(s-B)\[I=\frac{1}{2P(s-B)}\]